# Capacity Gains in MIMO Systems with Few-Bit ADCs Using Nonlinear Analog Circuits

### v2

@@ IV-B Communication Strategies and Achievable Rates
Given a threshold matrix $\mathbf{t}\!\in\!\mathbb{R}^{n_{q}\times{(\ell-1)}}$ and functions $f_{j}\in\mathcal{F}_{env}^{\delta}$, a quantizer $Q:\mathbb{R}\to[\ell]^{n_{q}}$ characterized by the tuple $(\ell,\delta,n_{q},f^{n_{q}},\mathbf{t})$ is defined as $Q(\cdot)\triangleq(Q_{1}(\cdot),Q_{2}(\cdot),\cdots,Q_{n_{q}}(\cdot))$, where $Q_{j}(y)\triangleq k$ if and only if $f_{j}(y)\in[t(j,k),t(j,k+1)],j\in[n_{q}]$. The associated partition of $Q(\cdot)$ is:
$\mathsf{P}=\{\mathcal{P}_{\mathbf{i}},\mathbf{i}\in[\ell]^{n_{q}}\}-\Phi,\text{ where }\mathcal{P}_{\mathbf{i}}=\{y\in\mathbb{R}|Q(y)=\mathbf{i}\},\mathbf{i}\in[\ell]^{n_{q}}.$
For a quantizer $Q(\cdot)$, we call $y\in\mathbb{R}$ a point of transition if the value of $Q(\cdot)$ changes at input $y$, i.e. if it is a point of discontinuity of $Q(\cdot)$. Let $r$ be a point of transition of $Q(\cdot)$. Then, there must exist output vectors $\mathbf{c}\neq\mathbf{c}^{\prime}$ and $\epsilon>0$ such that $Q(y)=\mathbf{c},y\in(r-\epsilon,r)$ and $Q(y)=\mathbf{c}^{\prime},y\in(r,r+\epsilon)$. So, there exists $j\in[n_{q}]$ and $k\in[\ell-1]$ such that $f_{j}(y)<t(j,k),y\in(r-\epsilon,r)$ and $f_{j}(y)\geq t(j,k),r\in(r,r+\epsilon)$, or vice versa; so that $r$ is a root of the function $f_{j,k}(y)\triangleq f_{j}(y)-t(j,k)$. Let $r_{1},r_{2},\cdots,r_{\gamma}$ be the sequence of roots of $f_{j,k}(\cdot),j\in[n_{q}],k\in[\ell-1]$ (including repeated roots), written in non-decreasing order, where $\gamma\triangleq(\ell-1)n_{q}2^{\delta}$. Let $\mathcal{C}=(\mathbf{c}_{0},\mathbf{c}_{1},\cdots,\mathbf{c}_{\gamma})$ be the corresponding quantizer outputs, i.e. $\mathbf{c}_{i-1}=\lim_{y\to r_{i}^{-}}Q(y),i\in[\gamma]$ and $\mathbf{c}_{\gamma}=\lim_{y\to\infty}Q(y)$. We call $\mathcal{C}$ the code associated with the quantizer and it plays an important role in the analysis provided in the sequel. Note that the associated code is an ordered set of vectors. The size of the code $|\mathcal{C}|$ is defined as the number of unique vectors in $\mathcal{C}$. Each $\mathbf{c}_{i}=(c_{i,1},c_{i,2},\cdots,c_{i,n_{q}}),i\in\{0,1,\cdots,\gamma\}$ is called a codeword. For a fixed $j\in[n_{q}]$, the transition count of position $j$ is the number of codeword indices where the value of the $j$th element changes, and it is denoted by $\kappa_{j}$, i.e., $\kappa_{j}\triangleq\sum_{k=1}^{\gamma}\mathbbm{1}(c_{i_{k}-1,j}\neq c_{i_{k},j})$. It is straightforward to see that $|\mathsf{P}|=|\mathcal{C}|$ since both cardinalities are equal to the number of unique outputs the quantizer produces. The following example clarifies the definitions given above.
Figure 6: The quantizer outputs in Example 2. The first four rows show the sign of the function $f_{j,k},j,k\in\{1,2\}$ for the values of $y$ within each interval. The last row shows the quantizer output in that interval.
@@
Let $n_{q}=\delta=2$ and $\ell=3$ and consider a quantizer characterized by analog processing functions $f_{1}(y)=A_{2}(y,(2,4))=||y-2|-4|$ and $f_{2}(y)=A_{2}(y,(4,0))=||y|-4|,y\in\mathbb{R}$, and thresholds
$t(1,1)=0,\quad t(1,2)=1,\quad t(2,1)=1,\quad t(2,2)=2,$
We have:
$f_{1,1}(y)=||y-2|-4|-1,\quad f_{1,2}(y)=||y-2|-4|$
$f_{2,1}(y)=||y|-4|-1,\quad f_{2,2}(y)=||y|-4|-2.$
The ordered root sequence is $(r_{1},r_{2},\cdots,r_{10})=(-6,-5,-3,$ 2,1,2,3,5,6,7)-2,-1,2,3,5,6,7)- 2 , - 1 , 2 , 3 , 5 , 6 , 7 ). The associated partition is:
$\mathsf{P}=\Big{\{}[-\infty,-6),(-6,-5),(-5,-3),(-3,-2),(-2,-1),(-1,2),$
@@
The associated code is given by $22,21,20,11,02,12,$ $11,10,01,12,22$. This is shown in Figure 6. The size of the code is $|\mathcal{C}|=8$. The high SNR capacity of a channel using this quantizer at the receiver is $\log{|\mathsf{P}|}=\log{|\mathcal{C}|}=\log{8}$.
IV-B1 Single Envelope Detector and One-bit Quantization
As a first step, we investigate scenarios with $\ell=2$ and $\delta=1$. We will build upon this to derive capacity expressions for $\delta,\ell\in\mathbb{N}$. It can be noted that since $\delta=1$, each $f_{j}(y)$ is of the form $|y-a_{j}|$ for some $a_{j}\in\mathbb{R}$. We sometimes write $f_{a_{j},j}(y)=|y-a_{j}|$ to explicitly denote the value of $a_{j}$. Given threshold $t_{j}>0$ the roots of $f_{a_{j},j}(y)-t_{j}$ are equal to $a_{j}\pm t_{j}$. The following proposition provides the high SNR capacity when $\ell=2,\delta=1$.
Proposition 1 (High SNR SISO Capacity with One-bit ADCs and Single-level Envelop Detectors).
Let $h\in\mathbb{R}$, $h\neq 0$, and $n_{q}>1$. Further assume that the ADC resolution is limited to $\ell=2$. Then,
$\lim_{P_{T}\to\infty}C_{env}(P_{T},P_{ADC},h,1)=1+\log{\lfloor\frac{P_{ADC}}{2\alpha}\rfloor},$
where $\alpha$ is the unit of power consumption for a one-bit ADC.
Proof.
Please see Appendix A. ∎
@@ IV-B1 Single Envelope Detector and One-bit Quantization
The following theorem provides a computable expression for the capacity under general assumptions on channel SNR.
Theorem 1 (SISO Capacity with One-bit ADCs and Single-level Envelope Detectors).
@@
Consider a system parametrized by $(P_{T},P_{ADC},h,\delta)$, where $P_{T},P_{ADC}>0,h\in\mathbb{R},n_{q}>1$, and $\delta=1,\ell=2$. Then, the capacity is given by:
$C_{env}(P_{T},P_{ADC},h,\delta)=\sup_{\mathbf{x}\in\mathbb{R}^{2n_{q}+1}}\sup_{P_{X}\in\mathcal{P}_{\mathbf{x}}(P_{T})}\sup_{\mathbf{t}\in\mathbb{R}^{2n_{q}}}I(X;\widehat{Y}),$ (2)
@@
where $n_{q}=\frac{P_{ADC}}{2\alpha}$, $\widehat{Y}=Q(hX+N)$, $N\sim\mathcal{N}(0,1)$, $\mathcal{P}_{\mathbf{x}}(P_{T})$ is the set of probability distributions defined on $\{x_{1},x_{2},\cdots,x_{2n_{q}+1}\}$ such that $\mathbb{E}(X^{2})\leq P_{T}$, and $Q(y)=k$ if $y\in[t_{k},t_{k+1}],k\in\{1,\cdots,2n_{q}]$ and $Q(y)=0$ if $y>t_{2n_{q}}$ or $y<t_{1}$.
Proof.
Please see Appendix B. ∎
IV-B2 Low-resolution ADCs and Concatenated Sequences of Envelope Detectors
Next, we consider systems with $\delta>1,\ell>2$. Recall that for $\delta>1$, each $f_{j}(y),j\in[n_{q}]$ is of the form $f_{j}(y)=A_{\delta}(y,a_{j}^{\delta})=|A_{\delta-1}(y,a_{j}^{\delta-1})-a_{j,\delta}|$, where $A_{1}(y,a_{j,1})=|y-a_{j,1}|$. For tractability, we use the notation $f_{a^{\delta}_{j},j}(y)\triangleq A_{\delta}(y,a_{j}^{\delta})$ to explicitly denote the bias vector $a^{\delta}_{j}$ used for the $j$th analog function. The following example introduces the concept of degenrate bias vectors for a given threshold matrix $\mathbf{t}$.
@@ IV-B2 Low-resolution ADCs and Concatenated Sequences of Envelope Detectors
Let $n_{q}=1,\ell=3,\delta=2$, and consider the thresholds $t_{1,1}=1$ and $t_{1,2}=2$. Given a bias vector, $(a_{1},a_{2})$, the associated analog function is $f_{a_{1}^{\delta},j}(y)=||y-a_{1}|-a_{2}|$. The ADC output is
$Q(y)=\begin{cases}0\qquad&\text{ if }||y-a_{1}|-a_{2}|<1\\ 1&\text{ if }1<||y-a_{1}|-a_{2}|<2\\ 2&\text{ if }2<||y-a_{1}|-a_{2}|\end{cases}.$
Note that if $a_{2}-1<0$, then this would be equivalent with:
$Q(y)=\begin{cases}0\qquad&\text{ if }|y-a_{1}|<1+a_{2}\\ 1&\text{ if }1<|y-a_{1}|<2+a_{2}\\ 2&\text{ if }2+a_{2}<|y-a_{1}|\end{cases}.$
@@
We call bias vectors $a_{j}^{\delta},j\in[n_{q}]$ which yield redundant envelope detectors, such as the one in Example 3, degenerate bias vectors. The following proposition characterizes the necessary and sufficient conditions for non-degeneracy of bias vectors.
@@
Let the threshold vector corresponding to the $j$th ADC be $t^{\ell-1}$, where $j\in[n_{q}]$. The bias vector of the corresponding analog operator $f_{j}(\cdot)$ is non-degenerate if and only if:
$0<t_{1}+\sum_{i=2}^{\delta}(-1)^{b_{i}}a_{i},\quad\forall b_{i}\in\{-1,1\}.$ (3)
The proof follows by noting that from definition $0<t_{1}<t_{2}<\cdots<t_{\ell-1}$ so that Equation (3) guarantees $0<t_{k}+\sum_{i=2}^{\delta}(-1)^{b_{i}}a_{i}$ for all $k\in[\ell-1]$, and is thus sufficient to ensure non-degeneracy. We will use the following notion of a fully-symmetric vector in deriving properties of roots of quantizers with non-degenerative bias vectors.
@@
A vector $\mathbf{b}=(b_{1},b_{2},\cdots,b_{2^{n}})$ is called symmetric if $b_{i}+b_{2^{n}-i}=b_{j}+b_{2^{n}-j},i,j\in[2^{n}-1]$. $\mathbf{b}$ is called fully-symmetric if it is symmetric and the vectors $(b_{1},b_{2},\cdots,b_{2^{n-1}})$ and $(b_{2^{n-1}+1},b_{2^{n-1}+2},\cdots,b_{2^{n}})$ are both fully-symmetric for $n>2$ and symmetric for $n=2$.
For instance, the vector $\mathbf{b}=(-7,-6,-5-4,4,5,6,7)$ is fully symmetric since it is symmetric and $(-7,-6,-5-4)$ and $(4,5,6,7)$ are both symmetric. As observed in previous section, the number and shape of the quantization Voronoi regions depends on the roots of the analog processing function. The following proposition characterizes useful properties of the roots which are used in the subsequent sections to derive capacity expressions of various communicaiton systems.
Proposition 3 (Properties of Roots of Associated Analog Functions).
Consider a quantizer $Q(\cdot)$ with threshold matrix $\mathbf{t}\in\mathbb{R}^{n_{q}\times(\ell-1)}$, and analog processing functions $f_{j}(\cdot),j\in n_{q}$, such that the corresponding bias vectors are non-degenerate and $f_{j,k}(\cdot)\triangleq f_{j}(\cdot)-t(j,k),j\in[n_{q}],k\in[\ell-1]$ do not have repeated roots. Let $r_{1},r_{2},\cdots,r_{\gamma}$ be the increasing sequence of roots, where $\gamma\triangleq(\ell-1)n_{q}2^{\delta}$. Then, there exists a partition $\{\mathcal{P}_{j,k},j\in[n_{q}],k\in[\ell-1]\}$ of $[\gamma]$ such that
@@
The proof follows by taking each $\mathcal{P}_{j,k}$ to be the ordered set of roots of $f_{j,k}$ for a given $j\in[n_{q}],k\in[\ell-1]$ and using properties of the absolute value. The following proposition states several useful properties for the code associated with a quantizer $Q(\cdot)$.
@@
Consider a quantizer $Q(\cdot)$ with threshold matrix $\mathbf{t}\in\mathbb{R}^{n_{q}\times(\ell-1)}$ such that $0<t(i,j)<t(i,j^{\prime}),i\in[n_{q}],j,j^{\prime}\in[\ell-1],j<j^{\prime}$, and analog processing functions $f_{j}(\cdot),j\in n_{q}$, such that the corresponding bias vectors are non-degenerate and $f_{j,k}(\cdot)\triangleq f_{j}(\cdot)-t(j,k),j\in[n_{q}],k\in[\ell-1]$ do not have repeated roots. The associated code $\mathcal{C}$ satisfies the following:
@@
    The number of codewords in $\mathcal{C}$ is equal to $\gamma\triangleq(\ell-1)n_{q}2^{\delta}+1$, i.e. $\mathcal{C}=(\mathbf{c}_{0},\mathbf{c}_{1},\cdots,\mathbf{c}_{\gamma-1})$.
@@
    All elements of the first codeword $\mathbf{c}_{0}$ are equal to $\ell-1$, i.e. $c_{i,0}=\ell-1,i\in\{0,1,\cdots,\gamma-1\}$.
@@
    Let $i_{1},i_{2},\cdots,i_{\kappa}$ be the non-decreasingly ordered indices of codewords, where the $j$th element has value-transitions. Then, the sequence $(c_{i_{1},j},c_{i_{2},j},\cdots,c_{i_{\kappa},j})$ is periodic, in each period it takes all values between 00 and $\ell-1$, and $|c_{i_{k},j}-c_{i_{k+1},j}|=1,k\in[\kappa-1]$ holds. Furthermore, $c_{i_{1},j}\in\{0,\ell-1\}$.

### v3

@@ IV-B The Quantizer and Its Associated Codebook
Given the ADC threshold matrix $t^{n_{q}\times{(\ell-1)}}\!\in\!\mathbb{R}^{n_{q}\times{(\ell-1)}}$ and implementable analog functions $f_{j}\in\mathcal{F}_{env}^{\delta},j\in[n_{q}]$, the quantizer $\mathbf{q}:\mathbb{R}\to[\ell]^{n_{q}}$ is defined as $\mathbf{q}(\cdot)\triangleq(q_{1}(\cdot)$ $,q_{2}(\cdot),\cdots,q_{n_{q}}(\cdot))$, where $q_{j}(y)\triangleq k$ if and only if [^3] For ease of notation, we have defined $t_{j,0}\triangleq-\infty$ and $t_{j,\ell}\triangleq\infty$. $f_{j}(y)\in[t_{j,k},t_{j,k+1}],j\in[n_{q}],k\in\{0,1,\cdots,\ell-1\}$. The associated partition is defined as:
$\mathsf{P}=\{\mathcal{P}_{\mathbf{c}},\mathbf{c}\in\{0,1,\cdots,\ell-1\}^{n_{q}}\}-\Phi,$ (7)
where $\mathcal{P}_{\mathbf{c}}=\{y\in\mathbb{R}|\mathbf{q}(y)=\mathbf{c}\},\mathbf{c}\in\{0,1,\cdots,\ell-1\}^{n_{q}}$.
For a quantizer $\mathbf{q}(\cdot)$, we call $y\in\mathbb{R}$ a point of transition if the value of $\mathbf{q}(\cdot)$ changes at input $y$, i.e. if it is a point of discontinuity of $\mathbf{q}(\cdot)$. Let $r$ be a point of transition of $\mathbf{q}(\cdot)$. There must exist output vectors $\mathbf{c}\neq\mathbf{c}^{\prime}$ and $\epsilon>0$ such that $\mathbf{q}(y)=\mathbf{c},y\in(r-\epsilon,r)$ and $\mathbf{q}(y)=\mathbf{c}^{\prime},y\in(r,r+\epsilon)$. So, there exists $j\in[n_{q}]$ and $k\in[\ell-1]$ such that $f_{j}(y)<t_{j,k},y\in(r-\epsilon,r)$ and $f_{j}(y)\geq t_{j,k},y\in(r,r+\epsilon)$, or vice versa; so that $r$ is a root of the function $f_{j,k}(y)\triangleq f_{j}(y)-t_{j,k}$. Note that there are a total of at most $(\ell-1)n_{q}2^{\delta}$ roots for the collection of functions $f_{j,k}(\cdot)$. Let $r_{1},r_{2},\cdots,r_{\gamma}$ be the sequence of roots of $f_{j,k}(\cdot),j\in[n_{q}],k\in[\ell-1]$ (including repeated roots), written in non-decreasing order, where $\gamma\leq(\ell-1)n_{q}2^{\delta}$. Let $\mathcal{C}=(\mathbf{c}_{0},\mathbf{c}_{1},\cdots,\mathbf{c}_{\gamma})$ be the corresponding quantizer outputs, i.e. $\mathbf{c}_{i-1}=\lim_{y\to r_{i}^{-}}\mathbf{q}(y),i\in[\gamma]$ and $\mathbf{c}_{\gamma}=\lim_{y\to\infty}\mathbf{q}(y)$. We call $\mathcal{C}$ the code associated with the quantizer. Note that the associated code is an ordered set of vectors. Each $\mathbf{c}_{i}=(c_{i,1},c_{i,2},\cdots,c_{i,n_{q}}),i\in\{0,1,\cdots,\gamma\}$ is called a codeword. The size of the code $|\mathcal{C}|$ is defined as the number of unique vectors in $\mathcal{C}$. For a fixed $j\in[n_{q}]$, the transition count of position $j$ is the number of codeword indices where the value of the $j$th element changes, and it is denoted by $\kappa_{j}$, i.e., $\kappa_{j}\triangleq\sum_{k=1}^{\gamma}\mathbbm{1}(c_{k-1,j}\neq c_{k,j})$. It is straightforward to see that $|\mathsf{P}|=|\mathcal{C}|$ since both cardinalities are equal to the number of unique outputs the quantizer produces. The following example clarifies the definitions given above.
Figure 4: The quantizer outputs in Examples 1 and 2. The first four rows show the sign of $f_{j,k},j\in[n_{q}],k\in[\ell-1]$ for $y$ within each interval, and the last row is the quantizer output.
@@
Consider the setting of Example 1. The functions $f_{j,k},j\in[n_{q}],k\in[\ell-1]$ are:
$f_{1,1}(y)=\big{|}|y-2|-3\big{|}-1,\qquad f_{1,2}(y)=\big{|}|y-2|-3\big{|}-2,$
$f_{2,1}(y)=\big{|}|y|-4\big{|}-1,\qquad f_{2,2}(y)=\big{|}|y|-4\big{|}-2.$
The ordered root sequence of the quantizer is $(r_{1},r_{2},\cdots,r_{13})=(-6,-5,-3,-2,0,1,2,3,4,5,6,7)$. The associated partition is:
$\mathsf{P}=\Big{\{}(-\infty,-6),[-6,-5),[-5,-3),[-3,-2),[-2,0),[0,1),[1,2)$
@@
The associated code is given by $22,21,20,11,02,12,$ $22,21,10,00,01,12,22$. This is shown in Figure 4. The size of the code is $|\mathcal{C}|=13$. The high SNR capacity of a channel using this quantizer at the receiver is $\log{|\mathsf{P}|}=\log{|\mathcal{C}|}=\log{13}$.
IV-C Single Envelope Detector and One-bit Quantization
As a first step, we investigate scenarios with one-bit ADCs and a single envelope detector per ADC, i.e., $\ell=2$ and $\delta=1$. Since $\delta=1$, each $f_{j,1}(y)$ is of the form $|y-b_{j}|$ for some $b_{j}\in\mathbb{R}$. Given an ADC threshold $t_{j,1}>0$ the roots of $f_{j,1}(y)-t_{j,1}$ are equal to $b_{j}\pm t_{j,1}$. So, there are a total of at most $2n_{q}$ roots for the quantizer. Then, the number of Voronoi regions of the quantization, which are segments of the real line separated by these roots, cannot be greater than $12n_{q}+$ 1. On the other hand, the first and last segments always have the same ADC outputs (for an example, refer to Figure 4). Consequently, the high SNR capacity is upper-bounded by $\log{2n_{q}}=\log{2\lfloor\frac{P_{ADC}}{2\alpha}\rfloor}$. The following proposition shows that this upper-bound is achievable. The proof is provided in [37].
Proposition 1.
Let $h\neq 0$, $P_{ADC}>0$, $\delta=1$, and $\ell=2$. Then,
$\lim_{P_{T}\to\infty}C_{env}(P_{T},P_{ADC},h,\delta)=1+\log{\lfloor\frac{P_{ADC}}{2\alpha}\rfloor},$ (8)
where $2\alpha$ is the unit of power consumption for a one-bit ADC.
The proof is provided in Appendix A.
To compare this result with the conventional approach, where analog processing is restricted to linear operators, note that without using envelope detectors and with $n_{q}$ one-bit ADCs one can partition the real line into at most $n_{q}+1$ Voronoi regions. Consequently, the capacity of such a system at high SNR is $\log{(n_{q}+1)}$. In comparison, when using envelope detectors the capacity is $\log{2n_{q}}$ which yields an improvement.
@@ IV-C Single Envelope Detector and One-bit Quantization
The following provides a computable expression for the capacity under general assumptions on SNR.
Theorem 1.
@@
Consider a system parametrized by $(P_{T},P_{ADC},h,\delta)$, where $P_{T},P_{ADC}>0,h\neq 0,n_{q}>1$, and $\delta=1,\ell=2$. Then, the capacity is given by:
$C_{env}(P_{T},P_{ADC},h,\delta)=\max_{\mathbf{x}\in\mathbb{R}^{2n_{q}+1}}\max_{P_{X}\in\mathcal{P}_{\mathbf{x}}(P_{T})}\max_{\mathbf{t}\in\mathbb{R}^{2n_{q}}}I(X;\widehat{Y}),$ (9)
@@
where $n_{q}=\lfloor\frac{P_{ADC}}{2\alpha}\rfloor$, $\widehat{Y}=q(hX+N)$, $N\sim\mathcal{N}(0,1)$, $\mathcal{P}_{\mathbf{x}}(P_{T})$ is the set of probability distributions defined on $\{x_{1},x_{2},\cdots,x_{2n_{q}+1}\}$ such that $\mathbb{E}(X^{2})\leq P_{T}$, and $q(y)=k$ if $y\in[t_{k},t_{k+1}],k\in[2n_{q}]$ and $q(y)=0$ if $y>t_{2n_{q}}$ or $y<t_{1}$.
The proof is provided in Appendix B. The theorem shows that, similar to the high-SNR regime, for low-SNRs the effect of using envelope detectors without concatenation is (almost) equivalent [^4] Note that the effect is not exactly equal to doubling the number of ADCs as the first and last quantization regions do not produce unique ADC outputs when using envelope detectors as can be observed in Figure 4. to increasing the number of ADCs from $n_{q}$ to $2n_{q}$.
IV-D Concatenated Envelope Detectors and Few-bit ADCs
Next, we consider systems with $\delta>1$ and $\ell>2$. Recall that for $\delta>1$, each $f_{j}(y),j\in[n_{q}]$ is of the form $f_{j}(y)=A_{\delta}(y,b_{j}^{\delta})=|A_{\delta-1}(y,b_{j}^{\delta-1})-b_{j,\delta}|$, where $A_{1}(y,b_{j,1})=|y-b_{j,1}|$. To evaluate the resulting channel capacity, we need to characterize the number and shape of Voronoi regions which can be produced using a quantizer consisting of concatenated envelope detectors and few-bit ADCs. To this end, we first introduce the notion of non-degenerate bias vectors (Proposition 2). Then, for envelope detectors with non-degenerate biases, we characterize some key properties of the associated code and the roots of the quantization function (Propositions 3 and 4). Subsequently, for any given sequence of roots satisfying the aforementioned properties, we provide a code construction mechanism which has that root sequence (Proposition 5). This enables us to derive the capacity expression for the communication systems under consideration (Theorem 3).
Some (degenerate) bias vectors lead to redundant envelope detectors. The following example clarifies the notion.
@@ IV-D Concatenated Envelope Detectors and Few-bit ADCs
Let $n_{q}=1,\ell=3,\delta=2$, and consider the thresholds $t_{1,1}=1$ and $t_{1,2}=2$. Given a bias vector $(b_{1},b_{2})$, the associated analog function is $f_{j}(y)=||y-b_{1}|-b_{2}|$. The ADC output is
$q(y)=\begin{cases}0\qquad&\text{ if }||y-b_{1}|-b_{2}|<1\\ 1&\text{ if }1\leq||y-b_{1}|-b_{2}|<2\\ 2&\text{ if }2\leq||y-b_{1}|-b_{2}|\end{cases}.$
Note that if $b_{2}-1<0$, then this would be equivalent with:
$q(y)=\begin{cases}0\qquad&\text{ if }|y-b_{1}|<1+b_{2}\\ 1&\text{ if }1+b_{2}\leq|y-b_{1}|<2+b_{2}\\ 2&\text{ if }2+b_{2}\leq|y-b_{1}|\end{cases}.$
@@
The following proposition characterizes the necessary and sufficient conditions for non-degeneracy of bias vectors.
@@
Let the threshold vector corresponding to the $j$th ADC be $t^{\ell-1}$, where $j\in[n_{q}]$. The bias vector $\mathbf{b}$ of the corresponding analog operator $f_{j}(\cdot)$ is non-degenerate if and only if:
$0<t_{1}+\sum_{i=2}^{\delta}(-1)^{p_{i}}b_{i},\quad\forall p_{i}\in\{-1,1\}.$ (10)
The proof follows by noting that from definition $0<t_{1}<t_{2}<\cdots<t_{\ell-1}$ so that Equation (10) guarantees $0<t_{k}+\sum_{i=2}^{\delta}(-1)^{p_{i}}b_{i}$ for all $k\in[\ell-1]$, and is thus sufficient to ensure non-degeneracy.
In the next step, we introduce the notion of a fully-symmetric vector which will be used in deriving properties of roots of quantizers with non-degenerative bias vectors.
@@
A vector $\mathbf{m}=(m_{1},m_{2},\cdots,m_{2^{n}})$ is called symmetric if $m_{i}+m_{2^{n}-i}=m_{j}+m_{2^{n}-j},i,j\in[2^{n}-1]$. $\mathbf{m}$ is called fully-symmetric if it is symmetric and the vectors $(m_{1},m_{2},\cdots,m_{2^{n-1}})$ and $(m_{2^{n-1}+1},m_{2^{n-1}+2},\cdots,m_{2^{n}})$ are both fully-symmetric for $n>2$ and symmetric for $n=2$.
For instance, the vector $\mathbf{m}=(-7,-6,-5-4,4,5,6,7)$ is fully symmetric since it is symmetric and $(-7,-6,-5-4)$ and $(4,5,6,7)$ are both symmetric. As observed in the proof of Theorem 1, a critical step in deriving the channel capacity resulting from a specific constellation of envelope detectors and ADCs is to characterize the number and shape of the Voronoi regions of the quantization operation resulting from the constellation. To this end, the following proposition provides useful properties of the roots which in turn determine the resulting Voronoi regions.
Proposition 3.
Consider a quantizer $\mathbf{q}(\cdot)$ with threshold matrix $t^{n_{q}\times(\ell-1)}\in\mathbb{R}^{n_{q}\times(\ell-1)}$, and analog processing functions $f_{j}(\cdot)\in\mathcal{F}_{env}^{\delta},j\in n_{q}$, such that the corresponding bias vectors are non-degenerate and $f_{j,k}(\cdot)\triangleq f_{j}(\cdot)-t_{j,k},j\in[n_{q}],k\in[\ell-1]$ do not have repeated roots. Let $r_{1},r_{2},\cdots,r_{\gamma}$ be the increasing sequence of roots, where $\gamma\triangleq(\ell-1)n_{q}2^{\delta}$. Then, there exists a partition $\{\mathcal{P}_{j,k},j\in[n_{q}],k\in[\ell-1]\}$ of $[\gamma]$ such that
@@
The proof follows by taking each $\mathcal{P}_{j,k}$ to be the ordered set of roots of $f_{j,k}(\cdot)$ for a given $j\in[n_{q}],k\in[\ell-1]$ and using properties of the absolute value operator. The following proposition states several useful properties for the code associated with a quantizer $\mathbf{q}(\cdot)$.
@@
Consider a quantizer $\mathbf{q}(\cdot)$ with threshold matrix ${t}^{n_{q}\times(\ell-1)}\in\mathbb{R}^{n_{q}\times(\ell-1)}$ such that $0<t_{i,j}<t_{i,j^{\prime}},i\in[n_{q}],j,j^{\prime}\in[\ell-1],j<j^{\prime}$, and analog processing functions $f_{j}(\cdot),j\in n_{q}$, such that the corresponding bias vectors are non-degenerate and $f_{j,k}(\cdot)\triangleq f_{j}(\cdot)-t_{j,k},j\in[n_{q}],k\in[\ell-1]$ do not have repeated roots. The associated code $\mathcal{C}$ satisfies the following:
@@
    The number of codewords in $\mathcal{C}$ is equal to $\gamma+1$, where $\gamma\triangleq(\ell-1)n_{q}2^{\delta}$, i.e. $\mathcal{C}=(\mathbf{c}_{0},\mathbf{c}_{1},\cdots,\mathbf{c}_{\gamma})$.
@@
    All elements of the first codeword $\mathbf{c}_{0}$ are equal to $\ell-1$, i.e. $c_{0,j}=\ell-1,j\in[n_{q}]$.
@@
    For $j\in[n_{q}]$, let $i_{1},i_{2},\cdots,i_{\kappa_{j}}$ be the non-decreasingly ordered indices of codewords, where the $j$th element has value-transitions. Then, the sequence $(c_{i_{1},j},c_{i_{2},j},\cdots,c_{i_{\kappa_{j}},j})$ is periodic, in each period it takes all values between 00 and $\ell-1$, and $|c_{i_{k},j}-c_{i_{k+1},j}|=1,k\in[\kappa_{j}-1]$ holds.